If an iron ray has a length of 100m when the temperature is 10°c by how much will it expand when the temperature rises to 30°c. Take linear expansivity of iron as 0.000012/degree celcius

Grade:12th pass

3 Answers

Meep
4 years ago
L2=L1(1+ coefficient of linear expansion(t2-t1)) 
=100*(1+0.000012(30-10))
=100.024
Thus, it expands by L2-L1 metres = 0.024m
 
 
Rajdeep
4 years ago
HELLO THERE!
Let’s collect the given data first.
Here’s what we are given:
 
Length of the iron rod (L) = 100m
Coefficient of linear expansion (\alpha) = 1.2\times 10^{-5}
Rise in temperature = 20 K
 
Now,
\triangle L = L \alpha \triangle T = 100 \times 1.2 \times 10^{-5} \times 20 \\= 2.4 \times 10^{-2} m
 
THANKS!
Bdya Poudel
3 years ago
Initial length of iron(L1)=100m
Initial temperature(T1)=10°C
Final omelette(T2)=40°C
Now,
Difference in length of iron between 40°C and 10°C=L1*alpha *∆T
        =100*1.13*10^-5*30
        =0.039
If there is the error of 0.039 in 100m,then in 2 Km I.e.2000m,there will be the error of=0.039*2000/100
    =0.78,,and//